# Bi-parental Reproduction may Enhance Species Sustainability by Conserving Shared Parental Traits more Faithfully than Mono-parental Reproduction

### v3

@@ 5 Visual Illustration by Simulation
Below we present the results of a few simulation run examples. These simulations are meant to serve as an accessible visual illustration for the general effects predicted by the above mathematical analysis; they were not subjected to elaborate quantitative analysis. In all examples we show here just one run; multiple runs using the same parameters and differing only by the pseudo-random numbers used at various steps in the reproduction process, yielded similar results.
@@ 5.1 Reproducing a single real number via a set of points
The first set of simulations follows directly the model described in Section 3 demonstrating the reproduction of a real number through an intermediate encoding by a cohort of points. The program source is available in [Marron et al., 2024] Figure 1 depicts the results. The orange color graphs refer to mono-parental reproduction (MR), and the blue to bi-parental reproduction (BR). The simulation was run with the following parameters.
@@
  • Initial phenotype is thus: $\sim 2.00$.
@@
We observe the following: (a) The average of the phenotypes of the entire cohort is closer to the original phenotype of approximately $2.00$ under BR than under MR. (b) The standard deviation within each cohort generation is smaller in BR than under MR; i.e., the similarity of members of a given generation to each other is higher under BR than under MR, and thus they are better positioned to replace each other within the ecosystem. (c) On average, the phenotypes of individuals of each generation, are closer to the average of the immediately preceding parent generation in BR than in MR. (d) On average, under BR, the phenotypes of individuals of each generation, are closer to the average of the original ancestral generation than under MR.
@@ 5.2 Reproducing images of printed text
The second set of simulations is inspired by the way the survival of ancient manuscripts over hundreds and thousands of years was enabled by reproduction, i.e., copying. Both “mono-parental” and “bi-parental” techniques were employed by scribes: most often copying from a single source, but sometimes collating multiple sources  [Skeat, 1999, pp.585;617]. (Note: we distinguish here between collation of several sources in the process of copying text, from the collation of sources by history researchers reconstructing an original document.) In this example we applied the reproduction techniques described above for the reproduction of images of text containing one word. We introduced a rudimentary measure of the preservation of properties as the success rate (across the cohort) of image recognition software (gpt-4-turbo model by OpenAI) in recognizing the text. This metric also hints at importance of traits in an organism’s interactions, which in this case can be seen as the sending and receiving of messages.
@@
Function F2. Noisy computation of a “phenotype”. Given $n$ copies of an image, as may be created using F1 above, compute the average of all respective pixel values; add noise to the result in the same manner as in F1, and resize the resulting image to the original resolution. This step may represent the collation process done by a person creating a fresh copy based on several sources, where each may have mistakes and defects.
The parameters used in the simulation are: seed image is an image of the word “TEACHING”, typed in the font Calibri Light at resolution 100x100 pixels; copies generated by F1 are 400x400 pixels; noise standard deviation is $h=20$; cohort size is $n=8$; and, the number of copies in the “image genotype” encoding is $k=8$. The simulation steps are as follows:
@@
    [Creating base cohort.] Repeat the following $n$ times:
@@
      Repeat $n$ times:
@@
Figure 2 depicts a sample of the resulting images, with one image from each generation.
Figure 3: The 8 individuals of generation 40 under BR and MR. See discussion in the body of the article about retention of common traits, both ancestral and emergent.
@@
We have also looked casually for common traits across the various cohorts and noticed that traits like contrast between the letters and their background, or the horizontal line of the letter T were preserved better under BR than under MR (see Figure 3). Under BR, we have also noticed conspicuous and sustained emergent traits, like a circle shape inside the letter C or a third vertical line in the letter H, where under MR we could not readily find any such emergent properties.
@@ 5.3 Reproducing images of flowers
Figure 4: Simulating reproduction of an image of a flower. See explanation in text.
In this illustration example we were inspired by object recognition in nature, as is done, for example, by insects species that pollinate only certain plant species. Figure 4 depicts the simulation of BR and MR of an image of a flower. The simulation uses the same process as in the case of text image, with the following parameters: The starting image is of an anemone flower, downloaded from https://it.pinterest.com/pin/red-anemones–797207571571931709/, and used under fair use license; the image was then cropped and resized to 100 x 100 pixels; noise standard variation is $h=10$; cohort size is $n=8$; Encoding set size is $k=8$; Number of generations is $80$; image recognition prompt is: “Your role is to identify the object in the image. Please provide a 1-word answer.";
One can see that, under BR, more of the images were recognized like the first one (“Poppy”) for more generations than under MR. We also observe that certain core properties, like being a flower were better preserved under BR.
Note: The images shown here are illustrative examples. While in all runs the image properties were sustained for longer under BR than under MR, some intermediate generations showed a local advantage for MR. Such results are to be expected, first, due to the random nature of the inserted noise, and perhaps more so, due to the possibility that in a particular cohort, an individual may reproduce several times, dominating the properties in the next generation.
@@ 6 Conclusion and Next Steps
We have shown that mathematically, bi-parental, or sexual, reproduction can contribute to conservation in the offspring of traits that are common and important in sustainment of the parent generation. Confirming the observation in nature, and investigating it in more elaborate models, including multiple properties, multiple generations, and more detailed reflection of biological aspects, remains as future work.

### v5

@@ 5 Visual Illustration by Simulation
Below we present the results of a few simulation run examples. These simulations are meant to serve as an accessible visual illustration for the general effects predicted by the above mathematical analysis; they were not subjected to elaborate quantitative analysis. For each simulation case we show here just one run; multiple runs using the same parameters and differing only by the pseudo-random numbers used at various steps in the reproduction process, yielded similar results.
@@ 5.1 Reproducing a single real number via a set of points
The first set of simulations follows directly the model described in Section 3 demonstrating the reproduction of a real number through an intermediate encoding by a cohort of points. The program source is available in [29]. Figure 1 depicts the results. The orange color graphs refer to mono-parental reproduction (MR), and the blue to bi-parental reproduction (BR). The simulation was run with the following parameters.
@@
  • Initial common phenotype is thus: $\sim 2.00$.
@@
We observe the following: (a) The average of the phenotypes of the entire cohort is closer to the original phenotype of approximately $2.00$ under BR than under MR. (b) On average, the phenotypes of individuals of each generation, are closer to the average of the immediately preceding parent generation in BR than in MR. (c) The standard deviation within each generation cohort is smaller under BR than under MR; i.e., the similarity of members of a given generation to each other is higher under BR than under MR. (d) On average, under BR, the phenotypes of individuals of each generation, are closer to the average of the original ancestral generation than under MR.
@@ 5.2 Reproducing images of printed text
The second set of simulations is inspired by the way the survival of ancient manuscripts over hundreds and thousands of years was enabled by reproduction, i.e., copying. Both “mono-parental” and “bi-parental” techniques were employed by scribes: most often copying from a single source, but sometimes collating multiple sources  [30](p.585;617). (Note: we distinguish here between collation of several sources in the process of copying text, from the collation of sources by history researchers reconstructing an original document.) In this example we applied the reproduction techniques described above for the reproduction of images of text containing one word. We introduced a rudimentary measure of the preservation of traits as the success rate (across the cohort) of image recognition software (gpt-4-turbo model by OpenAI) in recognizing the text. This metric also hints at the importance of traits in an organism’s interactions; in the present case, the interaction can be seen as the sending and receiving of messages.
@@
Function F2. Noisy computation of a “phenotype”. Given $k$ copies of an image, as may be created using F1 above, compute the average of all respective pixel values; add noise to the result in the same manner as in F1, and resize the resulting image to the original resolution. This step may represent the collation process done by a person creating a fresh copy based on several sources, where each may have mistakes and defects.
The parameters used in the simulation are: seed image is an image of the word “TEACHING”, typed in the font Calibri Light, black and white only, at resolution 100x100 pixels; copies generated by F1 are 400x400 pixels; noise standard deviation is $h=20$; cohort size is $n=8$; and, the number of copies in the “image genotype” encoding is $k=8$. The simulation steps are as follows:
@@
    [Creating base cohort.] Repeat the following $n=8$ times:
@@
      [ Create next generation.] Repeat $n=8$ times:
@@
Figure 2 depicts a sample of the resulting images, with one sample image from each sample generation.
@@
We have also looked casually for effects regarding additional common traits across the various cohorts. We noticed that traits like contrast between the letters and their background, or the horizontal line of the letter T were preserved better under BR than under MR (see Figure 4). Under BR, we have also noticed conspicuous and sustained emergent traits, like a circle shape inside the letter C or a third vertical line in the letter H, where under MR we could not readily find any such emergent traits.
@@ 5.3 Reproducing images of flowers
In this illustration example we were inspired by object recognition in nature, as is done, for example, by insects species that pollinate only certain plant species. Figure 3 depicts reproducing an image of a flower under BR and under MR. The simulation uses the same process as in the case of text image, with the following parameters: The starting image was one labeled as anemone flower, downloaded from (https://it.pinterest.com/pin/red-anemones--797207571571931709/), and used under fair use license; the image was then cropped and resized to 100 x 100 pixels; noise standard variation is $h=10$; cohort size is $n=8$; encoding set size is $k=8$; number of generations is $80$; image recognition prompt is: “Your role is to identify the object in the image. Please provide a 1-word answer.".
One can see that, under BR, images were assigned the same label as the first one, “Poppy”, for more generations than under MR. (We ignore the fact that the LLM’s first identification did match the label “anemone” originally assigned to the image.) We also observe that certain core traits, like being a flower were better preserved under BR.
Note: The images shown here are illustrative examples. While in all runs the image traits were sustained for longer under BR than under MR, some intermediate generations showed a local advantage for MR. Such results are to be expected, first, due to the random nature of the inserted noise, and perhaps more so, due to the possibility that in a particular cohort, an individual may reproduce several times, dominating the traits in the next generation.
@@ 6 Conclusion and Next Steps
We have shown that merging and averaging effects in bi-parental, or sexual, reproduction can contribute to conservation, in the offspring generation, of traits that are common in the parent generation, in a way that complements known biological processes like reproductive limitations and DNA repair. Confirming the observation in nature, and investigating it in more elaborate models, including multiple traits, ecosystem interactions, and incorporating more biological details, remains as future work.